A manufacturer of spark plugs knows that the process produces 4 percent defective. As part of
their quality assurance program, the manufacturer inspects 10 spark plugs each hour. If two or
more of the plugs are defective, the process is shut down and corrective action is taken.
a. Use the binomial distribution to determine the probability of finding two or more defective
plugs when the process is actually 4 percent defective.
b. Use the Poisson distribution to compute the same probability.

Answer by ewatrrr(24785)


Hi
a.  Using Binomial distribution: p(def) =.04   q = .96  n = 10
P(x ≥ 2) = 1 - P(x ≤ 1) = 1 - binomcdf(10, .04, 1) = 1 - .9418 = .0582

b.  Use the Poisson Distribution:  
P(x ≥ 2) = 1 - P(x ≤ 1) = 1 -poissoncdf(.04, 1) = 1-.9992 = .0008
Wish You the Best in your Studies.